undeadknight26 (undeadknight26):

Please help: The Martians ask you to explain one last thing, Ultimate Math Ambassador. They ask you to create a new function, h(x). Then assign any number to x. Using complete sentences, explain whether f(h(x)) and h(f(x)) will always result in the same number. You will use the function f(x) that you created in problem number 2.

undeadknight26 (undeadknight26):

The other function:f(x) = 7x + 2

undeadknight26 (undeadknight26):

H(x) = 2x - 4

undeadknight26 (undeadknight26):

H(5) = 2(5) – 4 H(5) = 10 – 4 H(5) = 6

undeadknight26 (undeadknight26):

@Idealist10

undeadknight26 (undeadknight26):

was this right? H(5) = 2(5) – 4 H(5) = 10 – 4 H(5) = 6 7(2(5) – 4) – 2 7(6) – 2 42 – 2 = 40

undeadknight26 (undeadknight26):

@asib1214 @Ashy98 @Whitemonsterbunny17 @RaphaelFilgueiras @Elsa213 @ganeshie8 @Hero

OpenStudy (idealist10):

Right now I don't have time, Please wait...

OpenStudy (idealist10):

H(5)=6 and f(5)=7*5+2=37.

undeadknight26 (undeadknight26):

I do not understand.

OpenStudy (idealist10):

f(h(x))=7(2x-4)+2 h(f(x))=2(7x+2)-4 ---------------- for example, if you plug in 5, f(h(5))=7(2*5-4)+2=7(10-4)+2=7*6+2=44 h(f(5))=2(7*5+2)-4=2(35+2)-4=2*37-4=70 ----------------------------------------- f(h(x)) and h(f(x)) will not result in the same number.

undeadknight26 (undeadknight26):

Oh i see. It wouldn't result with the same number because the base formula is too different from each other?

OpenStudy (anonymous):

isn't h(x)=f-1(x)?

undeadknight26 (undeadknight26):

Wouldn't that be an inverse equation then?

OpenStudy (anonymous):

h(x)=x^1/3, and f(x)=x^3

OpenStudy (anonymous):

the inverse has that property, but it's the unique solution?

undeadknight26 (undeadknight26):

I have no idea. I have submitted it already so i will wait and see.

OpenStudy (idealist10):

@RaphaelFilgueiras , but the functions aren't inverses.